Main INDEX
Monthly INDEX
PREV
NEXT
Make New Entry,
Make Followup Entry
User name efuchey
Log entry time 00:01:12 on March 2,2011
Entry number 349516
keyword=3/01/11 swing shift summary
Shift crew:
SL: Eric FUCHEY
TO: Whitney ARMSTRONG
3rd: Zhiwen ZHAO
4:50 pm: finishing recovery procedure, moved to LH2 4cm; Started DT studies;
5:10 pm: increasing current to 20 uA;
5:20 pm: going back to 15 uA, going to production -> 1M evts;
7:15 pm: production suspended for a while to try to fix dead time issues;
9:10 pm: going to controlled access;
9:30 pm: going back to beam permit, asked for 15 uA again;
9:50 pm: beam back !
10:45 pm: going to 30 uA;
11:10 pm: going back to 15 uA;
11:20 pm: going (again !) to controlled access;
11:35 pm: going back to beam permit;
11:55 pm: beam is back.
Run list:
2106: BeO, 5uA, Spot++;
2107: LH2 4cm, 15 uA, DT studies;
2108: LH2 4cm, 15 uA, DT studies;
2109: LH2 4cm, 15 uA, DT studies;
2110: LH2 4cm, 15 uA, DT studies;
2111: LH2 4cm, 15 uA, DT studies;
2112: LH2 4cm, 15 uA, DT studies;
2113: LH2 4cm, 15 uA, DT studies;
2114: LH2 4cm, 15 uA / 20 uA, DT studies;
2115: LH2 4cm, 15 uA, Production, short run;
2116: LH2 4cm, 15 uA, Production;
2117: LH2 4cm, 15 uA, Production;
2118: LH2 4cm, 15 uA, Production;
2119: LH2 4cm, 15 uA, Production;
2120: LH2 4cm, 15 uA, DT studies;
2121: LH2 4cm, 15 uA, DT studies;
2122: LH2 4cm, 15 uA, DT studies;
2123: LH2 4cm, 15 uA, DT studies;
2124: LH2 4cm, 15 uA, DT studies;
2125: LH2 4cm, 15 uA, Junk (coda problem);
2126: LH2 4cm, 15 uA, Junk (coda problem);
2126: LH2 4cm, 15 uA, Junk (coda problem);
2127: LH2 4cm, 15 uA, test;
2128: LH2 4cm, 15 uA, test;
2129: LH2 4cm, 15 uA, test;
2130: LH2 4cm, 15 uA, test;
2131: LH2 4cm, 15 uA, test;
2132: LH2 4cm, 15 uA, test;
2133: LH2 4cm, 15 uA, junk;
2134: LH2 4cm, 15 uA, test;
2135 -> 2139: LH2 4cm, 15 uA, junk;
2140: LH2 4cm, 15 uA, test;
2141: LH2 4cm, 30 uA, test;
2142: LH2 4cm, 30 uA, test;
2143: LH2 4cm, 15 uA, test;
2144: LH2 4cm, 15 uA, test;